OpenStudy (anonymous):

PLEASE HELP (LAST ONE) : A wooden railing is 82 m long. It is to be divided into four pieces. Three pieces will be the same length than each of the other three. Find the length of each piece!

OpenStudy (anonymous):

I'm not entirely sure you copied that problem properly.

OpenStudy (anonymous):

Lil B are you sure of the Q?

OpenStudy (anonymous):

OOPS!! I left out a line! Hold on! Sorry . . .

OpenStudy (anonymous):

A wooden railing is 82 m long. It is to be divided into four pieces. Three pieces will be the same length, and the fourth piece will be 2 m longer than each of the other three. Find the length of each piece!

OpenStudy (anonymous):

1st piece = x 2nd piece = x 3rd piece = x 4th piece = x + 2 Add them up x + x+ x + x + 2 = 82 add like terms 4x + 2 = 82 subtract 2 4x = 80 divide by 4 x = 20 The pieces are 20,20,20,22

OpenStudy (anonymous):

excellent ^_^!

OpenStudy (anonymous):

You should go back and read my post... It is great helping you, but you need to be able to do these on your own or you are going to have a really hard time....

OpenStudy (anonymous):

I have been doing them on my own. My problem is the actual equation part. I have all the components right, just in the wrong places or I have issues actually solving them.

OpenStudy (anonymous):

i think its also because I KNOW I need one-on-one attention and help when it comes to these! Again, I appreciate everyone's help and time! Good night!